Exercise 10-22 Interest capitalization [LO10-7]

On January 1, 2018, the Marjlee Company began construction of an office building to be used as its corporate headquarters. The building was completed early in 2019. Construction expenditures for 2018, which were incurred evenly throughout the year, totaled $5,400,000. Marjlee had the following debt obligations which were outstanding during all of 2018:

Construction loan, 12% $ 1,350,000
Long-term note, 11% 1,800,000
Long-term note, 8% 3,600,000


Required:
Calculate the amount of interest capitalized in 2018 for the building using the specific interest method.

Interest capitalized $648,000 selected answer incorrect

I need the correct answer- it is not 648000

Interest capitalized ?   
Answer #1

Answer = total capitalized interest = $283500

Explanation =

Average accumulated expenditure =

5400000/2 =2700000

Weighted average rate of all other debts:

1800000 × 11% 198000
3600000 × . 8% 288000
5400000 486000

486000/5400000 = 9%

Interest capitalized:

1350000 × 12% 162000
1350000(2700000-1350000) × 9% 121500
Total 283500